# The Fréchet Distance Unleashed: Approximating a Dog with a Frog

### v3

@@ 1.3.4 Result IV: Computing the Fréchet distance quickly for real inputs
The above leaves us with a natural strategy for computing the Fréchet distance between two given curves. Compute quickly, using simplification, a morphing between the two input curves, and maintain (using VE-Fréchet, for example) both upper and lower bounds on the true Fréchet distance. By carefully inspecting the morphing, (re)simplifying the curves in a way that is sensitive to their (local) Fréchet distance, and recomputing the above bounds, one can get an improved morphing. Repeat this process potentially several times till the upper and lower bounds meet, at which point the optimal Fréchet distance has been computed.
This seems somewhat overkill, but it enables us to compute (in practice) the exact Fréchet distance between huge polygonal curves quickly.
Figure 1.4: For these two curves, the solution involves an agent stopping at one point on one curve while the other agent traverses is zig-zag, and vice versa. The algorithm enforces monotonicity by refining the two curves by introducing new vertices. For the results, see here.
1.3.5 Main contribution: Implementation in Julia and Python
@@ 1.4 Significance
We demonstrate in this paper that a minor variant of the continuous Fréchet distance can be computed by a strikingly simple Dijkstra type algorithm. Furthermore, for many real world inputs it runs in near linear time (out of the box). Similarly, for many real world inputs the computed morphing is monotone, thus realizing the (standard) Fréchet distance. More importantly, because of the retractable nature of the morphing computed, the matching computed is more natural, and can handle noise/outliers more gracefully than the matchings computed by the current algorithms for the Fréchet distance. Indeed, areas that are noise/outliers, get isolated in the morphing to a small interval, that can be easily identified and handled, see Figure 1.2. We believe that this makes the morphings computed by our new algorithms significantly better for real world applications than previous algorithms.
@@
Finally, we implemented our new algorithms in Julia and Python, and made them publicly available as standard packages, thus making the computation of the Fréchet distance accessible to a wider audience. Using such packages in Python/Julia is significantly easier than using any not pre-compiled code in C++.
Summary.
The combination of simplification, new (and not so new) algorithmic ideas (such as retractablity via Dijkstra, and VE-Fréchet, among others), and careful implementation, leads to fast performance in practice beating other implementations. We discuss our implementation next — the algorithmic ideas described above are covered in detail in the appendices.
@@ 2 Implementation and experiments
Total RT in seconds
Dataset # pairs avg # points C++ Julia Julia MT C++/Julia
sigspatial $322,000$ $247.8$ $145$ $102$ $25$ $\mathbf{1.42}$
Characters $253,000$ $120.9$ $90$ $72$ $18$ $\mathbf{1.25}$
Geolife $322,000$ $1080.4$ $646$ $70$ $30$ $\mathbf{9.22}$
All tests combined 881 244 73 $\mathbf{3.61}$
Figure 2.1: Julia/C++ comparison, demonstrating that the Julia implementation is significantly faster. The C++ implementation is from Bringmann et al. bkn-wdfpa-21. Running times are in seconds. Julia MT stands for the multi-threaded implementation – since multi-threading is easy in Julia, and the task at hand is easily parallizable, we tested this, but of course this should not be taken as a direct (or remotely fair) comparison to the C++ implementation, and is provided for the reader amusement. In particular, the multi-threading done is pretty naive, and better performance should be possible by better fine-grained partition of the tasks. The tests were performed on a Linux system with 64GB memory with Intel i7-11700 CPU, with a decent 16 threads CPU, but far from the fastest hardware currently available. The “# pairs” column is the number of pairs of curves that had their Fréchet distance compared during this test. The “avg # points” is the average number of vertices per curve in this input set – underlying the benefit of simplification for the GeoLife input set.
@@ 2.1 Hardware and Implementations
Here, we discuss the individual implementations used in our experiments. The first two (Julia and Python) use the algorithmic techniques discussed in this work. The third is by Bringmann et al. bkn-wdfpa-21 written in C++, which uses completely different algorithmic ideas. We refer to each implementation by language for convenience, although our tasks are meant to evaluate effectiveness of the algorithmic approaches.
@@ 2.1.1 Julia
The Julia code is about 5000 lines for the library, and an additional 5000 for the examples/testing code. It implements the algorithms described here faithfully, with some additional hacks to handle floating point issues. Julia has (surprisingly easy to use) support for multi-threading (unlike C++ or Python), and we tested also a multi-threaded version (this is shown in the MT column in Figure 2.1). The anonymized code is available from https://frechet.xyz/, including instructions how to replicate our tests.
@@
We were able to significantly improve the performance of the Julia code using some standard low-level optimizations: (i) Julia performs inlining code automatically, but explicitly forcing it to inline some low-level key functions (such as computing the distance between two points), improved the performance measurably ($>10\%$). Julia also has high overhead for creating structs that get destroyed immediately (essentially because it allocates them on the heap, instead of the stack like C++). Thus, by rewriting the code so that computing the distance of a point from a segment is done directly on the endpoints of the segment (where the segment is an edge of a polygon) also improved performance. Caching simplifications, so that one can extract/compute new simplification quickly, among other ideas, see Section 5.6, were useful in creating fast code.
@@
Since Julia performance suffers when a lot of dynamic allocations are done (because of the automatic garbage collection), it turned out that implementing the retractable VE-Fréchet without using hashing (for the vertices of the graph) resulted in much faster code. To this end, one pre-allocates a quadratic size array representing the vertices of the grid (but exploring only the relevant portions of the grid/this array) — this works because the allocation/initialization of this (single!) array is extremely fast. This seems to yields significant speedups somewhere up to ten times faster than the hashing version, depending on the input (for large inputs, the two implementations seem to have similar running times).
@@ 2.1.2 Python
For the Python implementation, we represented curves using NumPy arrays numpy-20, where each row in the array is a single point on the input curve. This representation is memory efficient and allows for fast computations of distances between points using library primitives. We also made heavy use of Numba lps-nlpjc-15 to further improve performance, as many of the algorithms discussed in this work are iterative and thus easy to accelerate using the library.
@@
Although not exclusive to this implementation, the Python code suffers from some issues related to numerical precision. Specifically, this implementation struggles with underflow of floating point numbers, which seems to occur in situations where consecutive points are too close in the input data (i.e., the distance between points is too small). However, this doesn’t represent a meaningful distance between these points, and this can be resolved by adding simplification as a preprocessing step (such as the algorithm in Section 5.6).
Our experiments are meant to emphasize the performance of our algorithmic techniques, and we do not have any tasks specifically designed to evaluate issues with floating point numbers or methods which can be used to combat this problem. In particular, we did not encounter these issues in any significance with the datasets mentioned in this work. This is a potential avenue for future work if given data that encounters these issues frequently.
@@ 2.1.3 C++
The C++ implementation is provided by Bringmann et al. bkn-wdfpa-21, which seems to be the state of the art (they claim to be faster than any previously available implementation). The algorithm uses completely different algorithmic techniques to those of this work, providing a point of comparison for the scalability of our implementations.
Informally, the implementation of Bringmann et al. bkn-wdfpa-21 builds a decider for the Fréchet distance using a $kd$-tree over the free space diagram, keeping track of the reachable regions on the boundary of each cell, refining cells by continuing down the (virtual) $kd$-tree if needed.
Note, that unlike Bringmann et al. bkn-wdfpa-21, our implementation can compute the exact Fréchet distance — although in some cases, because of numerical issues the “exact” results are just (high quality) approximations, see Remark 1.6.
@@ 2.2 Experiments
To compare the implementations, we performed benchmarks using two different tasks using real-world datasets. These were selected to determine the scalability of the algorithms relative to the large size of real-world data. This section is organized by the different tasks performed.
@@ 2.2.1 Direct Fréchet distance computations
This was done on individual pairs of curves selected from different datasets. Specifically, we use curves taken from the GeoLife dataset zfxml-ggtdu-11 (as used in bkn-wdfpa-21), a stork migration dataset rktzw-eabgc-18, (retrieved from Movebank rktzw-deabg-18, used in bbknp-klmct-20), and a pigeon flight dataset pgrbvg-pigeon-17. Figure 2.2 lists the individual curves from each dataset used in our benchmarks.
We do not perform any type of transformations on the input data (to reduce aberration due to Earth’s curvature, for example), as we are primarily interested in the execution speed of our algorithm.
See Figure 2.4 and Figure 2.3 for the results (observe that the Julia code was tested for $1.001$ approximation – we have not done the same for the Python code, since it was numerically unstable in this case). The Python performance on these datasets is competitive with the Julia implementation, although in some cases significantly slower. In general, the Julia implementation is faster (we did not perform enough experiments to decide how much faster), but it is clear that the Python implementation is fast and robust enough to be used in practice. The runtimes also do not include the Numba JIT compilation times.
@@
2.2.2 Fréchet Decider
We compared our implementation performance in Julia to the C++ implementation. We evaluated our implementations with the same benchmarking task as Bringmann et al. bkn-wdfpa-21, using the provided code that downloads and generates Fréchet distance tests. We emphasize that this task and datasets were introduced prior to to this work, and were not selected to be favorable to the techniques introduced here.
This task used curves from three datasets; the dataset from the SIGSPATIAL GIS Cup wo-asgcr-2018, a dataset of handwritten characters (Characters) w-ct-08, and the GeoLife dataset zfxml-ggtdu-11. Each dataset is composed of numerous input curves, and then a list of pairs of curves together with a threshold, and the goal is to decide whether the Fréchet distance is above/below/equal to this threshold. See Figure 2.1. This task is naturally friendly to simplifications – we implemented a fast simplification data-structure. We iteratively simplify the two curves being compared to progressively finer resolutions, comparing them using the (monotonized) VE-Fréchet distance. If the comparison is outside the error interval, the result is returned, otherwise, the algorithm goes to the next finer resolution.
The GeoLife zfxml-ggtdu-11 dataset benefits tremendously from the simplification strategy, and the Julia implementation is more than $9$ times faster than the C++ code, see Figure 2.1. The worst dataset for the Julia implementation is the characters dataset w-ct-08 where it seems to be about $25\%$ faster than the C++ code. This dataset both does not benefit from simplification, and also seems to give rise to many non-monotone morphings. The Sigspatial seems be somewhat resistant to approximation, and our code is only about $40\%$ faster than the C++ code in this case. We emphasize that each test in Figure 2.1 involves the execution of well over $>250,000$ runs of the algorithm for different curves, so these results seems quite representative of real world performance.
As for the general C++ vs Julia performance comparison, this seems to be case dependent, with reported cases where either is faster. The common belief is that a careful enough implementation in C++ would be faster from Julia implementation (i.e., some of Julia libraries are implemented in C++).
In particular, the significant speed (646 seconds vs 70) improvement in the case of the GeoLife input suggests our algorithm is much faster in cases of longer curves where simplification really helps. Note that while our code in Julia uses a single thread, it uses an automatic garbage collector that runs in parallel threads.
@@ 2.3 Discussion
Overall, it is clear that our Julia implementation is faster than other currently available implementations for computing the Fréchet distance. Beyond that Julia seems like a great programming language to develop geometric algorithms — providing a combination of a safe, multi-threaded, high-level programming language with the performance of C++, without the pain involved in working in C++.
One optimization used by the Julia code, that should be generally useful, is the following. As a preprocessing step, precompute a hierarchy of simplified curves for each input curve. This improves the query process, but makes the preprocessing more expensive. Thus, storing such precomputed hierarchies might be a good idea if input curves are going to be used repeatedly for performing distance queries. We emphasize that the reported running times include this (light) preprocessing stage (interestingly, the SIGSPTIAL wo-asgcr-2018 competition allowed such preprocessing to not be included in the overall running time).
@@ 8 Pigeons C360_10 / C480_9
Figure 2.2: The inputs tested and where they are taken from, birds referring to the stork migration dataset rktzw-deabg-18, the GeoLife dataset zfxml-ggtdu-11, and Pigeons referring to a dataset from pgrbvg-pigeon-17.
@@
Figure 2.3: Julia performance on various inputs. All running times are is seconds. The ##\## columns specify the number of vertices in each input. The single missing running time is for a case where the program ran out of memory. The quality of approximation is in the title of the column. The VER column is for running the VE-Fréchet (retractable) algorithm on the original input curves, which is much slower than the exact algorithm, which computes the exact Fréchet distance, but uses simplification internally for speed.
@@ 3.1 The retractable path in a directed graph
Let $G=(V,E)$ be a directed graph with $n$ vertices and $m$ edges, with weights $\mathcalb{w}:E\rightarrow\mathbb{R}$ on the edges (assume for simplicity of exposition that they are all distinct). Consider a simple path $\pi$ in $G$. The bottleneck of $\pi$ is $b(\pi)=\max_{e\in\pi}\mathcalb{w}\left(e\right)$. For any two vertices $s$ and $t$ in $G$, let $\Pi(s,t)$ denote the set of all simple paths from $s$ to $t$ in $G$. The bottleneck distance between $s$ and $t$ is $d_{B}(s,t)=\min_{\pi\in\Pi(s,t)}b(\pi)$. The unique edge, (under our assumption of distinct weights on the edges), that realizes $d_{B}(s,T)$ is the bottleneck edge of $s$ and $t$, denote by $\mathcalb{b}(s,t)$.
@@
This variant of the problem is a $\min\max$, which is equivalent to the $\max\min$ (maximize cheapest edge on the path) via negation of the prices of edges. If the edges are sorted, then computing the bottleneck edge can be done in linear time. Dijkstra can be modified to solve this problem cktzz-bptdg-16, in $O(n\log n+m)$ time. Similarly, the problem is readily solvable in linear time if the underlying graph is a DAG. Gabow and Tarjan gt-atbop-88 showed an algorithm with running time $O(m\log^{*}m)$, and some minor further improvements are known cktzz-bptdg-16.
@@
For two vertices $s,t$ of a directed graph $G$ with distinct weights on its edges, the retractable path from $s$ to $t$, is the unique path that contains the edge $\mathcalb{b}(s,t)=u\rightarrow v$, and furthermore, the subpath from $s$ to $u$, and from $v$ to $t$ are both retractable.
Importantly, we are interested in computing the retractable path itself (not just the bottleneck). Confusingly, and this is not immediate to see, the standard dynamic programming algorithm for DAGs does not work in this case.
@@
It is not hard to see that a minor simplification to Dijkstra’s algorithm, started from $s$, leads to a $O(n\log n+m)$ time algorithm for computing the retractable tree, containing the retractable paths from $s$ to all the vertices in $G$ (we assume here that all of $G$ is reachable from $s$). Indeed, if $C$ is in the current set of vertices already visited, the algorithm always handles the next cheapest edge in the directed cut $(C,V\setminus C)$. As in Dijkstra’s algorithm, one can maintain a variable $d(v)$, to store the weight of cheapest edge from a vertex already visited to the (yet unvisited) vertex $v$. Setting $d(v)=\min(d(v),\mathcalb{w}\left(u\rightarrow v\right))$ when handling the edge $u\rightarrow v$, records this information. Under this vertex based accounting, the algorithm always next handles the vertex with minimum $d$ value that is not yet visited (and these values can be maintained in a heap). Gabow and Tarjan gt-atbop-88 described how to modify Dijkstra to solve the bottleneck problem – what we point in the next lemma is that it is somewhat stronger – it computes the retractable path.
@@
The running time follows as the algorithm performs $O(m)$ decrease-keys and $O(n)$ delete-min operations, which take $O(1)$ and $O(\log n)$ time respectively if using a Fibonacci heap.
As for correctness – we analyze a somewhat slower implementation of the algorithm (i.e., this is a variant of Prim’s algorithm using cuts). The algorithm repeatedly handles the cheapest edge not handled yet that is in the directed cut $(C,\overline{C})$333$(C,\overline{C})=\left\{u\rightarrow v\in E\left(G\right)\;\middle|\;u\in C,v\in V\left(G\right)\setminus C\right\}$., where $C$ is the set of vertices visited so far. Initially, $C=\{s\}$ and the min-heap is initialized to hold the set $(C,\overline{C})$. The algorithm now repeatedly extracts the minimum edge $u\rightarrow v$ in the heap. If $v\in C$, then the algorithm continues to the next iteration. Otherwise, the algorithm marks the $v$ as visited, and adds all the outgoing edges from $v$ to the heap. It is easy to verify (by induction) that this slower algorithm, and the algorithm described above compute the same bottleneck-tree.
As for the correctness of the slower algorithm – the proof is by induction on the number of edges of the retractable path. Consider the case that the retractable path is of length one – namely, the edge $e=s\rightarrow u$. Let $\mathcalb{w}=\mathcalb{w}\left(e\right)$. Clearly, before the algorithm handles $e$, all the weights being handled are strictly smaller than $e$, and thus the algorithm can not arrive to $u$. Since $e$ is scheduled by the algorithm, the claim readily follows.
So, consider a retractable path $\pi$ from $s$ to $u$ that contains more than one edge, and let $e=x\rightarrow y$ be its bottleneck edge. Let $\mathcalb{w}=\mathcalb{w}\left(e\right)$. Observe that the algorithm would handle all the edges of $\pi$ before handling any edge with strictly larger weight than $w$. As such, for our purposes, assume that $e$ is the heaviest edge in the graph. Let $S$ be all the vertices visited just before $e$ was handled, and let $T=V\setminus S$. Consider the two induced graphs $G_{S}$ and $G_{T}$, and observe that one can apply induction to both parts, separately, as there are no edges except $e$ between the two parts (there might be edges going “back”, but these can be ignored, as the vertices of $S$ were already handled by the algorithm). It thus follows that the algorithm computed $\pi$ as the desired path.

### v4

@@ 1.3.4 Result IV: Computing the Fréchet distance quickly for real inputs
The above leaves us with a natural strategy for computing the Fréchet distance between two given curves. Compute quickly, using simplification, a morphing between the two input curves, and maintain (using VE-Fréchet, for example) both upper and lower bounds on the actual Fréchet distance. By carefully inspecting the morphing, (re)simplifying the curves in a way that is sensitive to their (local) Fréchet distance, and recomputing the above bounds, one can get an improved morphing. Repeat this process potentially several times till the upper and lower bounds meet, at which point the optimal Fréchet distance has been computed.
This approach seems somewhat overkill, but it enables us to compute (in practice) the almost-exact Fréchet distance between huge polygonal curves quickly.
Figure 1.4: For these two curves, the solution involves an agent stopping at one point on one curve while the other agent traverses a zig-zag, and vice versa. The algorithm enforces monotonicity by refining the two curves by introducing new vertices. For the results, see here.
1.3.5 Contribution: Implementation in Julia and Python
@@ 1.4 Significance
We demonstrate in this paper that a minor variant of the continuous Fréchet distance can be computed by a straightforward Dijkstra-type algorithm. Furthermore, for many real-world inputs, it runs in near-linear time (out of the box). Similarly, for many real-world inputs, the computed morphing is monotone, thus realizing the (standard) Fréchet distance. More importantly, because of the retractable nature of the morphing computed, the matching computed is more natural, and can handle noise/outliers more gracefully than the matchings calculated by the current algorithms for the Fréchet distance. Indeed, areas that are noise/outliers get isolated in the morphing to a small interval, which can be easily identified and handled, see Figure 1.2. We believe that this makes the morphings computed by our new algorithms significantly better for real-world applications than previous algorithms. There is also previous work on other variants of Fréchet distance that are more robust to uncertainty and outliers [FNPR24], and shortcutting [DH13].
@@
Finally, we implemented our new algorithms in Julia and Python. We made them publicly available as standard packages, thus making the computation of the Fréchet distance accessible to a broader audience. Using such packages in Python/Julia is significantly easier than using any non-pre-compiled code in C++.
The combination of simplification, new (and not so new) algorithmic ideas (such as retractability via Dijkstra, and VE-Fréchet, among others), and careful implementation, leads to reasonably fast performance in practice.
1.5 Errata
An earlier version of our paper [HRR25] made the extremely incorrect claim that our Julia implementation is faster than the C++ implementation of Bringmann et al. [BKN21]. This error was due to both misconfiguration of the data and incorrect measurement of the running times of the C++ implementation. In reality, the C++ algorithm is, by several orders of magnitude, faster than our implementation/algorithm. We believe that the ideas in our paper may still be of interest. See [CHR25], who recently turned our approach into an exact algorithm after making some additional improvements, and provided an empirical evaluation.
@@ 2.1 Hardware and implementations
Here, we discuss the individual implementations used in our experiments. The first two (Julia and Python) use the algorithmic techniques discussed in this work. We emphasize, however, that our implementation is slower by several orders of magnitude than the implementation of Bringmann et al. [BKN21].
@@ 2.1.1 Julia
The Julia code is about 5000 lines for the library, and an additional 5000 for the examples/testing code. It implements the algorithms described here faithfully, with some additional hacks to handle floating-point issues. Julia has (surprisingly easy to use) support for multi-threading (unlike C++ or Python), and we also tested a multi-threaded version. The code is available from https://frechet.xyz/, including instructions on how to replicate our tests.
@@
We were able to significantly improve the performance of the Julia code using some standard low-level optimizations: (i) Julia performs inlining code automatically, but explicitly forcing it to inline some low-level key functions (such as computing the distance between two points), improved the performance measurably ($10\%$). Julia also has high overhead for creating structs that get destroyed immediately (essentially because it allocates them on the heap, instead of the stack like C++). Thus, by rewriting the code so that computing the distance of a point from a segment is done directly on the endpoints of the segment (where the segment is an edge of a polygon), performance was also improved. Caching simplifications, so that one can extract/compute new simplification quickly, among other ideas, see Section 5.6, helped create fast code.
@@
Since Julia performance suffers when a lot of dynamic allocations are done (because of the automatic garbage collection), it turned out that implementing the retractable VE-Fréchet without using hashing (for the vertices of the graph) resulted in much faster code. To this end, one pre-allocates a quadratic-sized array representing the vertices of the grid (but exploring only the relevant portions of the grid/this array) — this works because the allocation/initialization of this (single!) array is high-speed. This seems to yield significant speedups, somewhere up to ten times faster than the hashing version, depending on the input (for large inputs, the two implementations seem to have similar running times).
Julia mutable vs immutable behavior.
Another interesting technicality is the way Julia stores structs in arrays. If the struct is declared mutable (i.e., its fields can be changed), an array of such structs is implemented as an array of pointers to a singleton struct, each individually allocated on the memory heap. If struct is immutable, which is the default, such arrays are contiguous structs placed one after the other in memory (as in C++). Somewhat counterintuitively, an immutable struct can be modified by copying a newly constructed struct into it. This behavior gives rise to differences in performance. It is motivated by the idea that, say, points rarely require a single coordinate to be changed, and once created, they tend not to be modified. In any case, we changed our point type to be immutable, and it led to significant improvements in running time.
Parameter tuning.
A significant amount of speedup resulted from optimizing various parameters in the algorithm implementation. One can usually guess what should be roughly the right range for these parameters, but in the end, we manually searched over some values, which was both tedious and heuristic. The problem of parameter tuning is, of course, well known and widely studied in machine learning – and using such techniques might be the next stage in further optimizing the implementation.
@@ 2.1.2 Python
For the Python implementation, we represented curves using NumPy arrays [HMW+20], where each row in the array is a single point on the input curve. This representation is memory efficient and allows for fast computations of distances between points using library primitives. We also made heavy use of Numba [LPS15] to further improve performance, as many of the algorithms discussed in this work are iterative and thus easy to accelerate using the library.
@@
Although not exclusive to this implementation, the Python code suffers from some issues related to numerical precision. Specifically, this implementation struggles with underflow of floating-point numbers, which seems to occur in situations where consecutive points are too close in the input data (i.e., the distance between points is too small). However, this doesn’t represent a meaningful distance between these points, and this can be resolved by adding simplification as a preprocessing step (such as the algorithm in Section 5.6).
Our experiments are meant to emphasize the performance of our algorithmic techniques, and we do not have any tasks specifically designed to evaluate issues with floating-point numbers or methods that can be used to combat this problem. In particular, we did not encounter these issues to any significant extent with the datasets mentioned in this work. This is a potential avenue for future work if given data that encounters these issues frequently.
@@ 2.1.3 C++
The C++ implementation is provided by Bringmann et al. [BKN21], and it is faster by several orders of magnitude than our implementation.
@@ 2.2 Experiments
To compare our implementations, we performed benchmarks using two different tasks using real-world datasets. These were selected to determine the scalability of the algorithms relative to the large size of real-world data. This section is organized by the different tasks performed.
@@ 2.2.1 Direct Fréchet distance computations
This was done on individual pairs of curves selected from different datasets. Specifically, we use curves taken from the GeoLife dataset [ZFX+11] (as used in [BKN21]), a stork migration dataset [RKT+18a], (retrieved from Movebank [RKT+18], used in [BBK+20]), and a pigeon flight dataset [PGR+17]. Figure 2.1 lists the individual curves from each dataset used in our benchmarks.
We do not perform any transformations on the input data (to reduce aberration due to Earth’s curvature, for example), as we are primarily interested in the execution speed of our algorithm.
See Figure 2.3 and Figure 2.2 for the results (observe that the Julia code was tested for $1.001$ approximation – we have not done the same for the Python code, since it was numerically unstable in this case). The Python performance on these datasets is competitive with the Julia implementation, although in some cases significantly slower. In general, the Julia implementation is faster (we did not perform enough experiments to decide how much faster). Still, the Python implementation is fast and robust enough to be used in practice. The runtimes also do not include the Numba JIT compilation times.
@@ 2.3 Discussion
Overall, our Julia implementation is still significantly slower, by several orders of magnitude, than the C++ of Bringmann et al. [BKN21]. This leaves it open whether any of the ideas used in this paper can be used together with their algorithm to get a faster algorithm.
One optimization used by the Julia code, that should be generally useful, is the following. As a preprocessing step, precompute a hierarchy of simplified curves for each input curve. This improves the query process, but makes the preprocessing more expensive. Thus, storing such precomputed hierarchies might be a good idea if input curves are going to be used repeatedly for performing distance queries. We emphasize that the reported running times include this (light) preprocessing stage (interestingly, the SIGSPTIAL [WO18] competition allowed such preprocessing to not be included in the overall running time).
Acknowledgements.
We thank the authors of [CHR25] and [BKN21] for observing that our prior comparisons with the C++ implementation of Bringmann et al. [BKN21] were incorrect, as discussed in Section 1.5.
@@ 8 Pigeons C360_10 / C480_9
Figure 2.1: The inputs tested and where they are taken from, birds referring to the stork migration dataset [RKT+18], the GeoLife dataset [ZFX+11], and Pigeons referring to a dataset from [PGR+17].
@@
Figure 2.2: Julia performance on various inputs. All running times are in seconds. The #\# columns specify the number of vertices in each input. The single missing running time is for a case where the program ran out of memory. The quality of approximation is in the title of the column. The VER column is for running the VE-Fréchet (retractable) algorithm on the original input curves, which is much slower than the exact algorithm, which computes the exact Fréchet distance, but uses simplification internally for speed.
@@ 3.1 The retractable path in a directed graph
Let $G=(V,E)$ be a directed graph with $n$ vertices and $m$ edges, with weights $\mathcalb{w}:E\rightarrow\mathbb{R}$ on the edges (assume for simplicity of exposition that they are all distinct). Consider a simple path $\pi$ in $G$. The bottleneck of $\pi$ is $b(\pi)=\max_{e\in\pi}\mathcalb{w}\left(e\right)$. For any two vertices $s$ and $t$ in $G$, let $\Pi(s,t)$ denote the set of all simple paths from $s$ to $t$ in $G$. The bottleneck distance between $s$ and $t$ is $d_{B}(s,t)=\min_{\pi\in\Pi(s,t)}b(\pi)$. The unique edge (under our assumption of distinct weights on the edges) that realizes $d_{B}(s,T)$ is the bottleneck edge of $s$ and $t$, denoted by $\mathcalb{b}(s,t)$.
@@
This variant of the problem is a $\min\max$, which is equivalent to the $\max\min$ (maximize the cheapest edge on the path) via negation of the prices of edges. If the edges are sorted, then computing the bottleneck edge can be done in linear time. Dijkstra can be modified to solve this problem [CKT+16], in $O(n\log n+m)$ time. Similarly, the problem is readily solvable in linear time if the underlying graph is a DAG. Gabow and Tarjan [GT88] showed an algorithm with running time $O(m\log^{*}m)$, and some minor further improvements are known [CKT+16].
@@
For two vertices $s,t$ of a directed graph $G$ with distinct weights on its edges, the retractable path from $s$ to $t$, is the unique path that contains the edge $\mathcalb{b}(s,t)=u\rightarrow v$, and the subpath from $s$ to $u$, and from $v$ to $t$ are both retractable.
Importantly, we are interested in computing the retractable path itself (not just the bottleneck). Confusingly, and this is not obvious, the standard dynamic programming algorithm for DAGs does not work in this case.
@@
It is not hard to see that a minor simplification to Dijkstra’s algorithm, started from $s$, leads to a $O(n\log n+m)$ time algorithm for computing the retractable tree, containing the retractable paths from $s$ to all the vertices in $G$ (we assume here that all of $G$ is reachable from $s$). Indeed, if $C$ is in the current set of vertices already visited, the algorithm always handles the next cheapest edge in the directed cut $(C,V\setminus C)$. As in Dijkstra’s algorithm, one can maintain a variable $d(v)$ to store the weight of the cheapest edge from a vertex already visited to the (yet unvisited) vertex $v$. Setting $d(v)=\min(d(v),\mathcalb{w}\left(u\rightarrow v\right))$ when handling the edge $u\rightarrow v$, records this information. Under this vertex-based accounting, the algorithm always handles the next vertex with the minimum $d$ value that is not yet visited (and these values can be maintained in a heap). Gabow and Tarjan [GT88] described how to modify Dijkstra to solve the bottleneck problem – what we point out in the next lemma is that it is somewhat stronger – it computes the retractable path.
@@
The running time follows as the algorithm performs $O(m)$ decrease-keys and $O(n)$ delete-min operations, which take $O(1)$ and $O(\log n)$ time, respectively, if using a Fibonacci heap.
[…]
As for the correctness of the slower algorithm – the proof is by induction on the number of edges of the retractable path. Consider the case that the retractable path is of length one – namely, the edge $e=s\rightarrow u$. Let $\mathcalb{w}=\mathcalb{w}\left(e\right)$. Clearly, before the algorithm handles $e$, all the weights being handled are strictly smaller than $e$, and thus the algorithm can not arrive at $u$. Since the algorithm schedules $e$, the claim readily follows.
So, consider a retractable path $\pi$ from $s$ to $u$ that contains more than one edge, and let $e=x\rightarrow y$ be its bottleneck edge. Let $\mathcalb{w}=\mathcalb{w}\left(e\right)$. Observe that the algorithm would handle all the edges of $\pi$ before handling any edge with strictly larger weight than $w$. As such, for our purposes, assume that $e$ is the heaviest edge in the graph. Let $S$ be all the vertices visited just before $e$ was handled, and let $T=V\setminus S$. Consider the two induced graphs $G_{S}$ and $G_{T}$, and observe that one can apply induction to both parts, separately, as there are no edges except $e$ between the two parts (there might be edges going “back”, but these can be ignored, as the algorithm already handled the vertices of $S$). It thus follows that the algorithm computed $\pi$ as the desired path.